# Trapped-ion quantum simulation of the Fermi-Hubbard model as a lattice gauge theory using hardware-aware native gates

### v1

@@ VII QPU Results
Figure 3: Circuit executions on the Aria-1 QPU. Magnetization correlator $\chi_{ij}$ at five trotter steps of the trotterized $H_{LGT}$ built from $\hat{C}$ (in various shades of green, with label ”Explicit”) or $\hat{C}_{A}$ (in various shades of red, with label ”Ansatz”). $J=1.0$, $U=2.0$. Each circuit is subject the global spin conservation as a post selection error mitigation strategy. Each circuit goes through no compiler-level error mitigation, debiasing (D), or debiasing and sharpening (D + S). $\sigma_{ij}$ denotes the standard deviation in $\chi_{ij}$ at each trotter step across three circuit executions of each variant. Right: $\chi_{23}$ for $N=4$ sites, executed using 12 qubits. Left: $\chi_{34}$ for $N=6$ sites, executed using 18 qubits.
@@
Table 1: Average root mean square error (RMSE) of $\chi_{i,i+1}$ between noiseless and QPU simulation for five Trotter steps of each circuit type run three times. Right column denotes the average number of shots discarded from post-selection error mitigation by removing states that do not satisfy total spin conservation.
N=4 Circuits rmse % Shots discarded
Ansatz: Debiasing and Sharpening 0.079 52.07
Exact: Debiasing and Sharpening 0.106 52.31
Ansatz: Debiasing 0.097 51.81
Exact: Debiasing 0.108 52.41
Ansatz: No error mitigation 0.101 51.81
Exact: No error mitigation 0.124 51.65
N=6 Circuits rmse % Shots discarded
Ansatz: Debiasing and Sharpening 0.138 47.01
Exact: Debiasing and Sharpening 0.161 52.35
Ansatz: Debiasing 0.167 52.66
Exact: Debiasing 0.147 53.87
Ansatz: No error mitigation 0.165 43.18
Exact: No error mitigation 0.170 43.74
Ansatz: 10K shots 0.106 49.65
Exact: 10K shots 0.082 50.15
@@
This section presents our results from running the four-site and six-site Fermi-Hubbard model encoded using 12 and 18 qubits, respectively. We Trotterize the LGT Hamiltonian with equal time steps of $\Delta t=0.3$ for five Trotter steps each. The site qubits are initialized in a separable state such that each site in the original FHM has one fermion, with the left half of the lattice being populated with $\uparrow$ spins and the right half with $\downarrow$ spins, creating a domain wall in the middle of the lattice. The $N$ bond qubits are initialized as $\frac{1}{\sqrt{2}}(|+-\rangle^{\otimes(N/2)}+|-+\rangle^{\otimes(N/2)})$ for $N$ original fermion sites ($3N$ qubits). Due to cyclic boundary conditions, a domain wall also exists between site $i=1$ and site $i=N$. To benchmark circuit performance, we calculate quench dynamics using a two-point magnetization correlator $\chi_{ij}$ on opposite sides of the boundary between the up-spin domain and the down-spin domain (“domain wall”) in each case,
$\chi_{i,i+1}=\langle\hat{S}_{i}\hat{S}_{i+1}\rangle-\langle\hat{S}_{i}\hat{S}_{i+1}\rangle$ (22)
where $i=2$ for $N=4$ and $i=3$ for $N=6$, and $\hat{S}_{i}=\hat{n}_{i,\uparrow}-\hat{n}_{i,\downarrow}=\hat{S}_{i\uparrow}-\hat{S}_{i\downarrow}$.
@@
Each circuit is run in three ways: First without the error mitigation in Sec. VI.1, then with debiasing, and finally with both debiasing and sharpening. All circuits are subject to post-selection spin conservation, such that only states in the histogram that possess the correct number of up and down spins are retained and the histogram is renormalized.
@@
Figs. 3 summarizes the evolution of the $\chi_{i,i+1}$ at the domain wall for five Trotter steps with $\Delta t=0.3$, for the $N=4$ and $N=6$ site cases executed on the IonQ Aria-1 QPU. Debiasing consistently yields a lower variance over multiple circuit executions of the same circuit for each Trotter step, indicating consistent performance across multiple identical circuit preparations. Sharpening only improves circuit execution consistency for $N=6$, with more pronounced improvement for the explicit Trotter circuit compared to our native gateset ansatz reconstruction. Overall, we see comparable variance and increase in variance as Trotter step increases for the ansatz and explicit circuits, with debiasing consistently decreasing this variance.
@@
Table 1 summarizes the accuracy of each circuit’s $\chi_{i,i+1}$ evolution and the number of shots discarded as part of spin conservation post-selection. For the $N=4$ cases, we find an average of 23% lower root-mean-square-error (RMSE) for the ansatz compared to explicit evolution. For $N=6$, with the exception of evolution with debiasing, in which explicit evolution has a lower RMSE than the anstaz by 13%, the remaining two cases have an average of 9.8% lower RMSE from the ansatz. Debiasing consistently improved circuit performance, with a more profound decrease in RMSE when applied to the explicit circuits.
@@
Additionally, a prior work [13] has predicted that the number of shots per circuit run scales exponentially with $N$ for efficient post-selection. The result of this work is significantly more optimistic than this claim: Using one-way analysis of variance (ANOVA) at a significance level of $\alpha=0.05$, we find no statistically significant differences in the number of shots discarded due to conserved symmetry-based error mitigation for cases $N=6$ and $N=4$, regardless of other error mitigation strategies used, such as debiasing and sharpening. This is despite the fact that 1000 shots were used for each circuit execution regardless of circuit size or number of qubits. Additionally, when circuits were executed with 10,000 shots, the percentage of discarded shots remained comparable (Table 1).
VIII von Neumann entropy to further decrease two-qubit gatecount
In our above optimization, circuit depth, dictated by the number of layers ($l$), is a preselected hyperparameter. $\hat{C}_{A}$ with larger $l$ are more expressive , but comes at the cost of extra gates. Naturally we want to find $\hat{C}_{A}$ with minimal $l$ such that $\hat{C}\approx\hat{C}_{A}$. However, it is less efficient to conduct GBO at many different $l$ to determine the smallest $l$ and therefore the smallest gatecount that satisfies $\hat{C}\approx\hat{C}_{A}$. Below we outline an alternative approach.
Table 2: 2-qubit gate (2Q) cost per trotter step for modelling $N$ sites of the FHM encoded as a $\mathbb{Z}_{2}$. Results summarized for a) the initial explicit trotter step decomposition b) by parametrizing $\hat{C}$ and applying gradient-based optimization (GBO) on $\hat{C}_{A}$, and c) through applying GBO with the non Neumann Entropy (VNE) method on $\hat{C}_{A}$ and a parametrized $\hat{B_{A}}$. Our final method is a 36% reduction in 2Q cost per trotter step without comprising on exactness of the $\mathbb{Z}_{2}$ FHM encoding.
Trotter Step Variant 2Q Cost
Explicit 14N
GBO on $\hat{C}_{A}$ 12N
GBO and VNE on $\hat{C}_{A}$ and $\hat{B}_{A}$ 9N
VIII.1 Predicting the smallest depth of an ansatz
First, we append an additional XX gate between qubits $|S_{i,\sigma}\rangle$ and $|S_{i+1,\sigma}\rangle$ to create the ansatz in Fig 4. This ansatz ensures full single-qubit rotations for $l\geq 2$, and possible entanglement between each qubit in each layer. Therefore, the number and placement of entangling gates are the limiting factors to determine whether $\hat{C}_{A}$ can express $\hat{C}$.
$S(\rho_{i})=-Tr(\rho_{i}\log(\rho_{i}))$ (23)
Figure 4: a) Ansatz for the von Neumann entropy (VNE) optimization method. The ansatz number of layers is first truncated through the VNE method. Following this, gradient based optimization (GBO) is applied to find parameters such that $\hat{C}_{A}\approx\hat{C}$.During GBO, a number of single and two-qubit gate rotations go to $\mathcal{I}$, allowing further compression of the ansatz. b) Result of GBO over 13 parameters using IPG, Adam, and L-BFGS. Best cost history over three trials and 128 iterations for each optimizer. IPG achieves the highest fidelity circuit, approaching fidelity of $\mathcal{F}=1-10^{-7}$, with L-BFGS converging faster but to a lower fidelity (higher cost). c) Final pre-conditioner matrix $K$ from IPG, which estimates the inverse Hessian, similar to Fig. 2.
Here, $l\geq 3$ satisfied this expressibility criteria. Gradient-based optimization only converged to a solution at $l=4$, with the gates shown in Fig. 4 going to $\mathcal{I}$. The optimization sequence amounted to a thirteen 1Q gate reduction compared to $\hat{C}_{A}$, and more importantly, two fewer 2Q gates than $\hat{C}$. By also applying this process to $\hat{B}$, we were able to reduce $\hat{B}$ to an $XX(\theta)$ gate on $|\tau_{i-1,i}\rangle,|\tau_{i,i+1}\rangle$ and a single $R_{y}(\theta)$ on $|\tau_{i-1,j}\rangle$. Table 2 summarizes the overall reduction in two-qubit gates per Trotter step, with an overall reduction from 14N to 9N. This decrease is substantial: Assuming 400 2Q gate executions as the Aria-1 algorithmic qubit limit [69], we are able to increase the number of faithful Trotter steps simulated on the Aria-1 QPU by $>50\%$, from 7 to 11 for $N=4$, and from 4 to 7 for $N=6$.
@@ VIII.1 Predicting the smallest depth of an ansatz
Figure 5: Statevector simulations subject to a two-qubit depolarizing noise channel at different weights $\gamma$. Magnetization correlator $\chi_{ij}$ at eight trotter steps of the trotterized $H_{LGT}$ built from $\hat{C}$ (green) or $\hat{C}_{A}$ (red). $\hat{C}_{A}$ is constructed using the von Neumann Entropy (VNE) optimization method. $J=1.0$, $U=2.0$, and $\Delta_{t}=0.4$. 1000 shots for each circuit execution, except for 10,000 shot executions specified for the explicit Trotter circuit cases at $\gamma=0.005$. Error in $\chi_{ij}$ at each Trotter step is difference between each circuit and the noiseless result. Left: $\chi_{23}$ for $N=4$ sites. Right: $\chi_{34}$ for $N=6$ sites.
VIII.2 Benchmarking reduced circuit depth performance
Using $\hat{C}_{A}$ and $\hat{B_{A}}$ optimized via the method outlined in the previous section, we construct the overall Trotter step and compute the magnetization correlator across multiple trotter steps. This is achieved with a statevector simulator on a classical computer subject to a 2Q depolarizing noise channel defined as
@@ VIII.2 Benchmarking reduced circuit depth performance
Fig. 5 summarizes our results. For $N=4$ sites, we find an excellent representation of the dynamics of $\chi_{23}$ across all eight trotter steps by the ansatz constructed Trotter circuit for both $\gamma=0.001$ and $0.005$. This is in sharp contrast to the explicitly constructed Trotter circuit, which loses a faithful representation of $\chi_{23}$ past just two steps at $\gamma=0.005$. Although the explicitly constructed Trotter circuit ouputs a faithful evolution of $\chi_{23}$ at $\gamma=0.0001$, this is still at a higher error than the ansatz constructed Trotter circuit when subject to a noisier environment at $\gamma=0.005$.
For $N=6$ sites, we see a trend similar to the $N=4$ site case. The ansatz Trotter circuit maintains a low error, faithful representation of $\chi_{34}$ through all eight and 6 steps for $\gamma=0.001$ and $\gamma=0.005$ respectively. For the explicit Trotter circuit, we see faithful agreement across all eight steps at $\gamma=0.001$, but at significantly larger error compared to the ansatz trotter circuit. At $\gamma=0.005$, the explicit Trotter circuit loses faithful evolution of $\chi_{34}$ after four steps, two fewer than the ansatz pTrotter circuit.
For both $N=4$ and $N=6$ sites, executing the explicit Trotter circuits with 10,000 shots at $\gamma=0.005$ proved inconsistent in improving the explicit trotter circuit performance: For $N=6$, 10,000 shots resulted in lower error than 1000 shots, but still possessed a higher error than the ansatz Trotter circuits at $\gamma=0.005$. For $N=4$, no tangible improvement was observed.
IX Conclusion
In this work, we have demonstrated several strategies to substantially enhance gate-based simulation of a quantum many-body system modeled by a prototypical Fermi-Hubbard Hamiltonian while constraining the circuit depth. Strategies to mitigate the growing circuit size for an $N$-site lattice included the use of advanced optimization techniques such as IPG and the use of information theoretic metrics such as the von Neumann entropy to minimize the number of two-qubit entangling gates, resulting in an overall 36% reduction of 2-qubit gates. In addition, an ensemble of error mitigation strategies are applied to further enhance of the performance of simulation on current state-of-the-art quantum processing units (QPUs). These techniques include general-purpose procedures that come with the QPU execution such as sharpening and debiasing, as well as symmetry based post-selection that is specific to the model of interest, for example the global spin conservation and local charge conservation. The specific model we implemented is a $\mathbb{Z}_{2}$ mapping of the FHM on a trapped-ion quantum computer, whereby periodic boundary conditions were readily realized by harnessing all-to-all qubit connectivity. We show that a combination of these strategies allowed us to run deep quantum circuits at or beyond the gate count naively expected from gate infidelities.

### v2

@@ VI QPU and Noisy Statevector Simulator Results
Figure 4: a) Circuit executions on the Aria-1 QPU without von Neuman entropy (VNE) two-qubit gate optimization (see Fig. 5 for results with VNE optimization). Magnetization correlator $\chi_{34}$ at five Trotter steps for $N=6$ sites of the Trotterized $H_{LGT}$ built from $\hat{C}$ (in various shades of green, with label ”Direct”) or $\hat{C}_{A}$ (in various shades of red, with label ”Optimized”). Executed using 18 qubits. $J\Delta t=0.3$, $U\Delta t=0.6$. Each circuit is subject the global spin conservation as a post selection error mitigation strategy. Each circuit goes through no compiler-level error mitigation, debiasing (D), or debiasing and sharpening (D + S). b) Error denotes the absolute value difference between the ideal result and a circuit execution at each Trotter step. c) $\sigma_{34}$ denotes the standard deviation in $\chi_{34}$ at each Trotter step across three circuit executions of each variant.
@@
Table 2: Average root mean square error (RMSE) of $\chi_{i,i+1}$ between noiseless and QPU simulation for five Trotter steps of each circuit type run three times. Trotter step executions built from both the direct and optimized ansatz goes through either no compiler-level error mitigation strategy, debiasing (D), or debiasing and sharpening (D+S). Right column denotes the average number of shots discarded from post-selection error mitigation by removing states that do not satisfy total spin conservation.
Ansatz type RMSE % Shots discarded
Optimized: D + S 0.138 47.01
Direct: D + S 0.161 52.35
Optimized: D 0.167 52.66
Direct: D 0.147 53.87
Optimized: None 0.165 43.18
Direct: None 0.170 43.74
This section presents our results from running the six-site Fermi-Hubbard model encoded using 18 qubits.
VI.1 Simulation parameters
@@
We Trotterize the LGT Hamiltonian with equal time steps. Site qubits are initialized in a separable state such that each site in the original FHM has one fermion, with the left half of the lattice being populated with $\uparrow$ spins and the right half with $\downarrow$ spins, creating a domain wall in the middle of the lattice. The $N$ bond qubits are initialized as $\frac{1}{\sqrt{2}}(|+-\rangle^{\otimes(N/2)}+|-+\rangle^{\otimes(N/2)})$ for $N$ original fermion sites ($3N$ qubits). Due to cyclic boundary conditions, a domain wall also exists between site $i=1$ and site $i=N$. To benchmark circuit performance, we calculate quench dynamics using a two-point magnetization correlator $\chi_{ij}$ on opposite sides of the boundary between the up-spin domain and the down-spin domain (“domain wall”) in each case,
$\chi_{i,i+1}=\langle\hat{S}_{i}\hat{S}_{i+1}\rangle-\langle\hat{S}_{i}\rangle\langle\hat{S}_{i+1}\rangle$ (19)
where $i=3$ for $N=6$, and $\hat{S}_{i}=\hat{n}_{i,\uparrow}-\hat{n}_{i,\downarrow}$
@@
Circuits executed on the IonQ Aria-1 QPU are run in three ways: First without the error mitigation, then with debiasing, and finally with both debiasing and sharpening. All circuits are subject to post-selection spin conservation, such that only states in the histogram that possess the correct number of up and down spins are retained and the histogram is renormalized.
VI.2 QPU Results
@@
For circuits executed on a QPU, subcircuits are optimized without the additional VNE optimization outlined in Sec. IV.4. Figs. 4 summarizes the evolution of the $\chi_{i,i+1}$ at the domain wall for five Trotter steps with $\Delta t=0.3$, for $N=6$ sites on the IonQ Aria-1 QPU. Debiasing consistently yields a lower variance over multiple circuit executions of the same circuit for each Trotter step, indicating consistent performance across multiple identical circuit preparations. Sharpening improvess circuit execution consistency, with more pronounced improvement for the direct Trotter circuit compared to our optimized native gateset ansatz reconstruction. Overall, we see comparable variance and increase in variance as Trotter step increases for both the optimized and direct ansatz, with debiasing consistently decreasing this variance.
@@
Table 2 summarizes the accuracy of each circuit’s $\chi_{i,i+1}$ evolution and the number of shots discarded as part of spin conservation post-selection. With the exception of evolution with debiasing, in which the directly implemented Trotter steps have a lower root-mean-square-error (RMSE) than the anstaz by 13%, the remaining two cases have an average of 9.8% lower RMSE from the optimized ansatz. Debiasing consistently improved circuit performance, with a more profound decrease in RMSE when applied to the direct circuits.
@@
Additionally, a prior work [13] has predicted that the number of shots per circuit run scales exponentially with $N$ for efficient post-selection. The result of this work is significantly more optimistic than this claim: Using a one-way analysis of variance (ANOVA) test at a significance level of $\alpha=0.05$, we find no statistically significant difference in the percentage of shots discarded due to conserved symmetry-based error mitigation. This makes the number of shots discarded independent of whether debiasing and sharpening are used, despite debiasing and sharpening requiring an order-of-magnitude greater number of shots.
VI.3 Noisy statevector results
@@
Figure 5: a) Circuit executions on a classical statevector simulator with von Neuman entropy (VNE) 2-qubit gate optimization. $\chi_{34}$ for $N=6$ sites. Statevector simulator subject to a 2-qubit depolarizing noise channel at different weights $\gamma$. Magnetization correlator $\chi_{ij}$ at eight trotter steps of the trotterized $H_{LGT}$ built from $\hat{C}$ (green) or $\hat{C}_{A}$ (red). $J\Delta t=0.4$, $U\Delta t=0.8$. 1000 shots for every circuit execution. b) Error in $\chi_{ij}$ at each Trotter step is difference between each circuit and the noiseless result.
Subcircuits for our noisy statevector results were optimized using the VNE method outlined in Sec. IV.4. The statevector simulator is subject to a two-qubit depolarizing noise channel defined as
@@
Fig. 5 summarizes our results. We find an excellent representation of the dynamics of $\chi_{34}$ across seven trotter steps by the Trotter circuit constructed by the optimized ansatz for both $\gamma=0.001$ and $0.005$. This is in sharp contrast to the directly constructed Trotter circuit, which loses a faithful representation of $\chi_{34}$ past just two steps at $\gamma=0.005$ and four steps at $\gamma=0.001$.
VII Conclusion
In this work, we have presented an algorithm-hardware co-design methodology for executing deep, Trotterized digital quantum simulations of strongly-correlated systems on digital quantum computers. Our methodology is the integration of four core components: First, symmetry-constrained three-qubit sub-circuits that are parametrized and classically optimized via iteratively pre-condition gradient descent (IPG); second, expressibility guided two-qubit gate-count reduction using the von Neumann Entropy (VNE) of each sub-circuit; third, general purpose compiler-level error mitigation via debiasing and sharpening; and fourth, post-selection based error mitigation specific to the symmetry of the model of interest.
The specific instance we implemented is a six-site $\mathbb{Z}_{2}\$LGT mapping of the Fermi-Hubbard Model (FHM) on a trapped-ion quantum computer, whereby periodic boundary conditions were readily realized by harnessing all-to-all qubit conenctivity. Our metholody yielded substantial circuit compilation gains: IPG optimization on a three-qubit subcircuits alone reduced the native two-qubit gate count per Trotter step by 14%. This, combined with compiler-level debiasing and symmetry-based post-selection from global spin conservation error mitigation allowed for a faithful observation of domain-wall quench dynamics even at total circuit fidelities between $10^{-1}$ to $10^{-4}$. VNE compression yielded an additional 12% two-qubit gate count reduction, bringing the total reduction to 36%. This doubled the number of Trotter steps that could be simulated under a depolarizing noise channel on a clasical statevector simulator.